Amanda is decorating her home and wants to buy some house plants. She is interested in three types of plants costing $7, $10, and $13. If she has budgeted $150 for the plants and whats to buy exactly 15 of them, what are her options?
Answer by lwsshak3(11628)

Amanda is decorating her home and wants to buy some house plants. She is interested in three types of plants costing $7, $10, and $13. If she has budgeted $150 for the plants and whats to buy exactly 15 of them, what are her options?
let x=number of $7 plants
let y=number of $10 plants
let z=number of $13 plants
..
by trial and error:
x=5, y=5, z=5
x+y+z=15
5*7+5*10*5*13=35+50+65=$150
one of her options: buy 5 of each type of plant